OpenStudy (anonymous):

Need help step by step please! csc^2x/1+tan^2x

jimthompson5910 (jim_thompson5910):

hint: 1 + tan^2 = sec^2

OpenStudy (anonymous):

I got that far and it chives me csc^2x/sec^2x but from there I'm stomped

jimthompson5910 (jim_thompson5910):

that's the same as \[\Large \left(\frac{\csc(x)}{\sec(x)}\right)^2\]

jimthompson5910 (jim_thompson5910):

what does \[\Large \frac{\csc(x)}{\sec(x)}\] turn into ?

OpenStudy (michele_laino):

please note that: \[\csc x = \frac{1}{{\sin x}}\]